OpenStudy (anonymous):

i dont understand these types of problems can someone help: The lengths of 2 sides of the triangle are given. The length of the third side must be greater than ____ but less than __. some examples: 6,9 15,13 7n,10n

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

if the sides of the triangle are a,b,c and if you know the sides of a,b, then the third side c is subject to the following a-b < c < a+b

jimthompson5910 (jim_thompson5910):

in this case, a > b

OpenStudy (anonymous):

yeah i get that part

jimthompson5910 (jim_thompson5910):

for example, say we have a triangle with sides of 10, 12, x in this case, a = 12, b = 10, c = x so... a-b < c < a+b 12-10 < x < 12+10 2 < x < 22 so the third side is between 2 and 22 (not including 2 and not including 22)

OpenStudy (anonymous):

yeah i get that you add them and subtract them but for one the numbers are 15 and 20 and the answer key says 3 and 15

jimthompson5910 (jim_thompson5910):

well if the two given sides are known to be 15 and 20, then the third side c is between 20-15 = 5 and 20+15 = 35 so 5 < c < 35

jimthompson5910 (jim_thompson5910):

I have no idea how they got 3 and 15. Three is too small. Also, you can go higher than 15.

OpenStudy (anonymous):

oh i was reading the wrong one it was for 6 and 9

jimthompson5910 (jim_thompson5910):

so 9-6 = 3 and 9+6 = 15 which I'm sure you see by now

OpenStudy (anonymous):

yeah i got it now thanks so much for the help!!

jimthompson5910 (jim_thompson5910):

you're welcome